OpenStudy (anonymous):

7w+4 = 4w+3 ----- ----- +5 6 7

jigglypuff314 (jigglypuff314):

multiply both sides by the least common multiple of 6 and 7 which is 42 so you would get 7(7w+4) = 6(4w+3) + 42(5) then distribute 49w+21 = 24w+18+210 then simplify 49w+21 = 24x+ 228 do you know how to go about solving it now?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

25w=207 w=8.28?

jigglypuff314 (jigglypuff314):

yep :)